What is the total time of flight of an object projected vertically towards with a speed of 30 ms-
1
Expert's answer
2021-08-18T02:24:18-0400

The time consists of time up and time down:


"t=t_u+t_d,\\\space\\
t_u=t_d=\frac vg,\\\space\\
t=2\frac vg=6.12\text { s}."